OpenStudy (anonymous):

find the equation of the tangent line to f(x)=e^(x^2-x) at x=0

OpenStudy (mathmale):

1) Find the derivative of f(x)=e^(x^2-x). This represents the slope of the T. L. 2) Substitute 0 for x in this derivative. The result is the slope of the T. L. at x=0. What next?

OpenStudy (s3a):

Steps: 1) Find f'(x). 2) Plug in x = 0 into f'(x). 1) f'(x) = e^(x^2 - x) * (x^2 - x)' f'(x) = e^(x^2 - x) * (2x - 1) 2) f'(0) = e^(0^2 - 0) * (2*0 - 1) f'(0) = e^0 * (-1) f'(0) = -1